OpenStudy (anonymous):

What is the solution to the system of equations? x+y= 5 x-2y= 2 (___,___) PLZ HELP I WILL MEDAL AND FAN!!!

OpenStudy (compassionate):

Do you know the two methods for solving systems?

OpenStudy (compassionate):

Well, substituting and eliminating. This method looks like it'll be easier to substitute. x+y= 5 x-2y= 2 So in your first equation x + y = 5, solve for y. y = 5 - x Now in your second eqauation, plug 5 - x into y to get x.

OpenStudy (anonymous):

@livias.random

OpenStudy (anonymous):

x+y= 5 x-2y= 2 Rearrange: x + y = 5 y = 5 - x Substitute: x - 3y = 2 x - 2 (5 - x) = 2 x - 10 + 2x = 2 3x - 10 = 2 3x = 12

OpenStudy (anonymous):

3x = 12 x = 4 x + y = 5 4 + y = 5 y = 1

OpenStudy (anonymous):

ok i am writing this down. I am following

OpenStudy (anonymous):

ok so is it (1,12) or (12,1)

OpenStudy (anonymous):

Well; x = 4 y = 1 Typically written as (4,1)

OpenStudy (anonymous):

oh ok i have two more like this can u help me?